a feather of mass 50g and a stone of mass 75g fall from the top of a cliff at the same time. if the height of the cliff is 32m, calculate the time taken by the stone to reach the ground neglecting air resistance. (g = 10m/s2)
1
Expert's answer
2019-06-11T10:00:40-0400

in case of no air resistance

"h = \frac{gt^{2}}{2},"

respectively

"t =(\frac{2h}{g})^{0.5} = 6.4^{0.5} = 2.5 (s)."